OpenStudy (anonymous):

(x-3)^2=4(11-x)? what is it? i need help PRONTO.

OpenStudy (anonymous):

x=7 or -5

OpenStudy (anonymous):

(x-3)^2=4(11-x) x^2-6x+9=44-4x x^2-2x-35=0 (x-7)(x+5)=0 So x=7 or x=-5

OpenStudy (anonymous):

(a-b)^2=a^2-2ab+b^2